# Stochastic Sampling of Operator Growth Dynamics

### v1

@@ Abstract
In Fig. 3d, we plot $\Delta r_{n}$, and remarkably, we find an intriguing two-plateau structure. We infer that the spectral density features a crossover between two exponential tails. Indeed, if we suppose
$\Phi(\omega)\sim\begin{cases}e^{-|\omega|/\omega_{1}}&|\omega|\lesssim\omega_{c}\\ e^{-|\omega|/\omega_{2}}&|\omega|\gtrsim\omega_{c}\end{cases}$ (12)
@@
where $\omega_{1}$ and $\omega_{2}$ are the decay rates and $\omega_{c}$ is the crossover scale. Using a saddle point approximation of Eq. 4, we may find that
$\Delta r_{n}\approx\begin{cases}2\omega_{1}/e&n\lesssim n_{1}=\omega_{c}/2\omega_{1}\\ 2\omega_{2}/e&n\gtrsim n_{2}=\omega_{c}/2\omega_{2}\end{cases}$ (13)
@@
In particular, we have $n_{1}\omega_{1}=n_{2}\omega_{2}=\omega_{c}/2$. To show our numerical results are consistent with Eq. 12, we plot $n\Delta r_{n}$ in the inset of Fig. 3d, and find a remarkably large $\omega_{c}\approx 90J$. This crossover would have been missed in a brute-force calculation.
@@
Owing to the thermal nature of the dynamics, we expect the large $n$ behavior of $r_{n}$ to be universal for typical local operators with a given set of model parameters. In [43], we indeed see this anticipated behavior for boundary $X$ operator, with $\log(n)\Delta r_{n}\simeq 2.2(1)J$ and $\Delta r_{n}\simeq 1.26(4)J$ in 1D and 2D respectively, consistent with the plots here. We also find that the structure of the intermediate plateau as seen in 2D is not universal. Nevertheless, the asymptotic plateau value is reached only at a similar large value $n_{2}\approx 40$. Correspondingly, the high-frequency tail of the spectral density is truly established only at a frequency that was numerically inaccessible before the present work.
Statistical mechanics perspective of operator growth – From the point of view of statistical mechanics, it is natural to examine the geometric properties of the operator string configurations Fig. 1b. In particular, the configurations with large $n_{\max}$ describe the vicinity of the critical point $\tau\nearrow\tau_{*}$ in the partition function (3) (in 1D $\tau_{*}=\infty$ because of the log-correction). Therefore, these large $n_{\max}$ configurations exhibit universal, scale-invariant behavior ubiquitous to continuous phase transitions.
@@
Figure 4: Scaling collapse of the average operator support size (for the $Z$ operator), as function of $k$, for different values of $n_{\max}$, in 1D (a), where $S_{\max}:=n_{\max}/\log n_{\max}$, and in 2D (b).
@@
To illustrate this point, we measured the average support size of the operator string as a function of temporal index $k$ in configurations of an increasing range of maximal expansion order $n_{\max}$. The results, plotted in Fig. 4, are in agreement with the following scaling ansatz in the large $n_{\max}$ limit:
@@
where $f_{\text{2D}}$ and $f_{\text{1D}}$ are scaling functions, which capture the evolution of the operator size growth to an $n_{\max}$-dependent maximum (and then decrease) in terms of rescaled time $k/2n_{\max}$ and size $S/S_{\max}$.
The maximal size $S_{\max}\sim n_{\max}$ in 2D, and we expect the same in higher dimensions as well, since this is the only asymptotics compatible with the universal moment growth in  Eq. 5. The slow convergence to the scaling collapse echos the late establishment of the universal growth observed above. In 1D, the maximal size $S_{\max}=n_{\max}/\log(n_{\max})$ has a log-correction, which has the same geometric origin as the log-correction to the moment growth: in 1D, the operator can only grow at the two extremities, which is a severe entropy penalty. Therefore the dominant contributions must be a compromise: the operator grows less in order to “scramble” more in the bulk. Previous arguments [44, 36] indicated a three-stage process where the operator grows to $\sim S_{\max}$ at $k\sim S_{\max}$, then stays as such until $2n_{\max}-k\sim S_{\max}$ at which point the operator starts to shrink. In units of rescaled time $k/2n_{\max}$, the growth and shrinking stages’ duration would be vanishing $\sim 1/\log(n_{\max})$, and the scaling function $f_{\text{1D}}$ would be a constant. In Fig. 4 we see indeed that a plateau establishes itself and extends to the whole interval as $n_{\max}$ increases, however very slowly (due to the log dependence). In 2D, one may worry that a similar “boundary-versus-bulk” entropy penalty applies as well. However, delicate rigorous arguments confirm that this is not the case [36]. Put simply, these arguments state that the operator support in 2D can have a “rough” boundary (with length comparable to the bulk area) and thus enough ways to grow in order to generate the moment asymptotics (5). Our results on moment ratio asymptotics provide strong (yet indirect) evidence of this unexpected geometric fact, whose direct observation will require pushing the numerics to large $n$.
@@
Discussion – We conclude our presentation by flagging several research directions motivated by our results. Our approach can be extended to models whose associated Liouvillian dynamics are expressed as sign problem-free sums. The quantum XY model is a notable example [37], which crucially admits a global $\text{U}(1)$ symmetry and hence can provide access to transport properties of conserved charges [47, 48]. The presence of disorder is also compatible with our method, it will be interesting to investigate the signature of Many-Body Localization in the evolution of operator strings. Another important direction is extracting real frequency dynamical spectral functions away from the high-frequency limit, as can, in principle, be via Lanczos coefficients expressed in terms of moment ratios [49, 50, 51]. We note that the latter might suffer from high sensitivity to statistical noise and may require an improvement on current algorithms. Lastly, we expect the support size scaling functions to be universal and valid for a generic class of chaotic Hamiltonians, and predicting them theoretically is an interesting open question. We leave the aforementioned possibilities to future efforts.
@@
Acknowledgements – We thank Ehud Altman and Assa Auerbach for fruitful discussions. S.G. acknowledges support from the Israel Science Foundation (ISF) Grant no. 586/22 and the US–Israel Binational Science Foundation (BSF) Grant no. 2020. U.B. acknowledges support from the Israel Academy of Sciences and Humanities through the Excellence Fellowship for International Postdoctoral Researchers. Computational resources were provided by the Intel Labs Academic Compute Environment and the Fritz Haber Center for Molecular Dynamics, Hebrew University.
Supplemental Materials:“Stochastic Sampling of Operator Growth Dynamics”
I The Monte Carlo algorithm
In this section, we outline our Marko chain Monte Carlo (MCMC) sampling algorithm. We first derive the precise form of the partition function taking into account the re-weighting procedure. Following this, we describe each of the Monte Carlo moves and derive their transition probabilities based on the detailed balance conditions.
We start by deriving Eq. 8 in the paper. Combining $\mathcal{Z}(\tau=1/J)$ from Eq. 3 with Eq. 6 in the main text, the partition function becomes:
$\mathcal{Z}(\tau=1/J)=\sum_{n=0}^{\infty}\sum_{\mathcal{S}_{2n}}\frac{1}{J^{2n}(2n)!}\prod_{k=0}^{2n-1}(O_{k+1}|\mathcal{L}_{i_{k},a_{k}}|O_{k})$ (S1)
where we recall that $\mathcal{S}_{2n}=\{(i_{k},a_{k})\}_{k=0}^{2n-1}$ is a sequence of position-type pairs that specify the local Liouvillians $\mathcal{L}_{i_{k},a_{k}}$. To proceed, we truncate this partition sum to sequence $n\geq n_{\max}$. We also fix the Liouvillian sequences to be of length $2n_{\max}$, by inserting $2n_{\max}-2n$ factors of the identity Liouvillian $\mathcal{L}_{0,0}$ in all possible combinations. This procedure however generates duplicate terms which are compensated for by dividing with the combinatorial factor (${2n_{\max}}\choose{2n_{\max}-2n}$. This gives us the desired expression of the truncated partition sum:
$G=\sum_{\mathcal{S}_{2n_{\max}}}\frac{1}{J^{2n}(2n)!}\times\frac{(2n_{\max}-2n)!(2n)!}{(2n_{\max})!}\prod_{k=0}^{2n_{\max}-1}(O_{k+1}|\mathcal{L}_{i_{k},a_{k}}|O_{k}).$ (S2)
As mentioned in the main text, the expansion order distribution associated with $G$ falls typically exponentially on either side of its modal point. In practice, this hinders the sampling of all the $n_{\max}+1$ sectors in an ergodic way. To flatten out this distribution, we employ a variant of the Wang-Landau algorithm [40, 42] that implements a re-weighting procedure. Before explaining our scheme, it is illuminating to rewrite $G$ using a set of generalized weights $\{w_{n}\}$:
$G^{\prime}=\sum_{\mathcal{S}_{2n_{\max}}}w_{n}\frac{(2n_{\max}-2n)!(2n)!}{(2n_{\max})!}\prod_{k=0}^{2n_{\max}-1}(O_{k+1}|\mathcal{L}_{i_{k},a_{k}}|O_{k})\,=\sum_{n}w_{n}\mu_{2n}.$ (S3)
@@ I The Monte Carlo algorithm
The weights $\{w_{n}\}$ are initialized as $w_{n}=1/(J^{2n}(2n)!)$. In practice one needs to only keep track of weight ratios between consecutive $n$ sectors $w_{n}/w_{n+1}$. The reweighting procedure works iteratively. At each step, we take $N$ samples and determine the number of times $c_{n}$ that each sector is visited by the MCMC. We define $c_{n}$ as:
$c_{n}=\sum_{i=1}^{N}\delta_{n,n_{i}}\,.$ (S4)
@@
With generalized weights, we need to rescale the counts in Eq. 9 with $\{w_{n}\}$ to obtain the moment ratios. Our observables now become:
$r_{n}^{2}=\frac{\mu_{2n+2}}{\mu_{2n}}=\frac{\left\langle\delta_{n^{\prime},n+1}\right\rangle_{n^{\prime}}}{\left\langle\delta_{n^{\prime},n}\right\rangle_{n^{\prime}}}\times\frac{w_{n}}{w_{n+1}}.$ (S5)
@@
I.1 Add Drop
Figure S1: Illustration of the add-drop moves with (a) $\mathcal{L}_{Z}$ Liouvillians (b) $\mathcal{L}_{X}$ Liouvillians (c) $\mathcal{L}_{ZZ}$ Liouvillians. The temporal direction is vertical.
The add-drop move is designed to allow transitioning between the different $n$ sectors by adding and removing a pair of identical local Liouvillians. As illustrated in Fig. S1, the move is defined as follows:
@@ I.1 Add Drop
We now determine the transition probabilities. Let $i$ be a configuration in the $n-1$ sector with the $(k-1)$-th and $k$-th Liouvillian slots unoccupied. Let $f$ be a configuration in the $n$ sector that is identical to $i$ except with these slots occupied at site $r$ by Liouvillians $\mathcal{L}_{r_{k-1},a_{k-1}}=\mathcal{L}_{r_{k},a_{k}}=:\mathcal{L}_{r,a}$. The two configurations are thus related by an add-drop move. Based on Eq. S3, detailed balance imposes the constraint:
$\frac{T_{i\rightarrow f}A_{i\rightarrow f}}{T_{f\rightarrow i}A_{f\rightarrow i}}=\frac{P_{f}}{P_{i}}=\frac{w_{n}(O_{k-1}|\mathcal{L}_{r,a}|O_{k})(O_{k}|\mathcal{L}_{r,a}|O_{k+1})(2n)(2n-1)}{w_{n-1}(2n_{\max}-2n+2)(2n_{\max}-2n+1)}\,.$ (S6)
Here, $P_{i},P_{f}$ are the respective weights, $T_{i\rightarrow f},T_{f\rightarrow i}$ are probabilities to propose the add/drop transitions and $A_{i\rightarrow f},A_{f\rightarrow i}$ are the acceptance probabilities for the moves. Now, proposing an add operation involves choosing a time slice, followed by choosing a site and then choosing a Liouvillian type; proposing a drop amounts to simply choosing a time slice. Therefore the proposition probabilities are
$T_{i\rightarrow f}=\frac{1}{2n_{\max}+1}\times\frac{1}{L^{d}}\times\frac{(O_{k-1}|\mathcal{L}_{r,a}|O_{k})(O_{k}|\mathcal{L}_{r,a}|O_{k+1})}{\sigma}\,,\,T_{f\rightarrow i}=\frac{1}{2n_{\max}+1}\,.$ (S7)
@@
$A_{i\rightarrow f}=\text{min}\left(1,\frac{w_{n}L^{d}\sigma(2n)(2n-1)}{w_{n-1}(2n_{\max}-2n+2)(2n_{\max}-2n+1)}\right)\,,\,$ (S8)
$A_{f\rightarrow i}=\text{min}\left(1,\frac{w_{n-1}(2n_{\max}-2n+2)(2n_{\max}-2n+1)}{w_{n}L^{d}\sigma(2n)(2n-1)}\right)\,.$ (S9)
I.2 Swap
Figure S2: Illustration of the swap move. We show two examples of Liouvillians that commute with each other: (a) $\mathcal{L}_{Z}$ and $\mathcal{L}_{ZZ}$ Liouvillians sharing a site. (b) Two $\mathcal{L}_{ZZ}$ Liouvillians sharing just one site. In addition to (a) and (b), identical Liouvillians (those of the same kind, on the same site; and along the same axis in case of $\mathcal{L}_{ZZ}$) and those on different sites with no overlap, trivially commute.
The swap move permutes through different configurations in the same $n$ sector using the commutation property of Liouvillian pairs, see Fig. Fig. S2 for an illustration. Before discussing the steps, we first define the notion of “ceiling” and “floor”. The ceiling (floor) of a Liouvillian is the temporally closest Liouvillian above (below, respectively) it which doesn’t commute with it. If a non-commuting Liouvillian is lacking, the ceiling/floor is then defined to be the initial/final temporal boundary. Now, the swap steps are as follows:

### v2

@@ V Numerical Results
In Fig. 4d, we plot $\Delta r_{n}$ vs $n$. We observe that $\Delta r_{n}$ settles into an intermediate false plateau $\Delta r_{n}\simeq 1.6J$ in the range $12\lesssim n\lesssim 20$ before transitioning to its true asymptotic plateau of $\Delta r_{n}\simeq 1.24J$ that is truly established beyond $n\approx 40$. Since a constant value of $\Delta r_{n}$ corresponds to an exponentially decaying spectral function with decay rate $\omega_{0}=\Delta r_{n}/2$, the two-plateau structure allows us to infer that the spectral density features a crossover between two exponential tails. If we suppose
$\Phi_{ZZ}(\omega)\sim\begin{cases}e^{-|\omega|/\omega_{1}}&|\omega|\lesssim\omega_{c}\\ e^{-|\omega|/\omega_{2}}&|\omega|\gtrsim\omega_{c}\end{cases}$ (12)
@@
where $\omega_{1}$ and $\omega_{2}$ are the decay rates and $\omega_{c}$ is the crossover scale. Using a saddle point approximation on Eq. 4, we find that
$\Delta r_{n}\approx\begin{cases}2\omega_{1}&n\lesssim n_{1}=\omega_{c}/2\omega_{1}\\ 2\omega_{2}&n\gtrsim n_{2}=\omega_{c}/2\omega_{2}\end{cases}$ (13)
@@
In particular, we have $n_{1}\omega_{1}\approx n_{2}\omega_{2}\approx\omega_{c}/2$. To show that our numerical results are consistent with Eq. 12, we plot $n\Delta r_{n}$ in the inset of Fig. 4d, and find $n_{1}\approx 20$ and $n_{2}\approx 30$. This gives a remarkably large $\omega_{c}=2n_{1}\omega_{1}\approx 32J$. We note that the brute force calculations, limited to the first $16$ moments, only capture the intermediate spectral decay rate $\omega_{1}$, missing the true asymptotic decay rate $\omega_{2}$ entirely.
@@
Owing to the thermal nature of the dynamics, we expect the large $n$ behavior of $r_{n}$ to be universal for typical local operators with a given set of model parameters. In Appendix B, we indeed see this anticipated behavior for boundary $X$ operator, with $\log(2n)\Delta r_{n}\simeq 2.2(1)J$ and $\Delta r_{n}\simeq 1.26(4)J$ in 1D and 2D respectively, consistent with the plots here. We also find that the structure of the intermediate plateau as seen in 2D is not universal. Nevertheless, the asymptotic plateau value is reached only at a similar large value $n_{2}\approx 40$. Correspondingly, the high-frequency tail of the spectral density is truly established only at a frequency that was numerically inaccessible before the present work.
VI Statistical mechanics perspective of operator growth
@@
Figure 5: Scaling collapse of the average operator support size (for the $Z$ operator), as function of $k$, in 1D (a), where $S_{\max}:=n_{\max}/\log n_{\max}$, and in 2D (b).
From the point of view of statistical mechanics, it is natural to examine geometric properties of the operator string configurations Fig. 1b. In particular, configurations with large $n_{\max}$ describe the vicinity of the critical point $\tau\nearrow\tau_{*}$ in the expansion Eq. 3 (in 1D $\tau_{*}=\infty$ because of the log-correction). Therefore, these large $n_{\max}$ configurations exhibit universal, scale-invariant behavior ubiquitous to continuous phase transitions.
@@
To illustrate this point, we measured the average support size of operator strings against temporal index $k$ in configurations with an increasing range of maximal expansion order $n_{\max}$. The results, plotted in Fig. 5, agree with the following scaling ansatz in the large $n_{\max}$ limit:
@@
where $f_{\text{2D}}$ and $f_{\text{1D}}$ are scaling functions, which capture the evolution of the operator size growth to an $n_{\max}$-dependent maximum (and then decrease) in terms of rescaled time $k/2n_{\max}$ and size $S/S_{\max}$. The scaling of $S_{\max}$ with respect to $n_{\max}$ is expected to be universal, and depend on the dimension. In 2D and higher dimensions, $S_{\max}\sim n_{\max}$. This is the only asymptotics compatible with the universal moment growth in  Eq. 5. Indeed, the slow convergence to the scaling collapse echos the late establishment of the universal growth observed above. In 1D, there is a log-correction $S_{\max}=n_{\max}/\log(n_{\max})$. This has the same geometric origin as the log-correction to the moment growth: the operator can only grow at the two extremities, which is a severe entropy penalty. Therefore the dominant contributions are a compromise: the operator grows less in order to “scramble” more in the bulk, and $S_{\max}$ is the optimal scaling form of the maximal size [45, 38]. Finally, the scaling functions $f_{\text{2D}}$ and $f_{\text{1D}}$ are independent of the operator, which corresponds to the boundary condition of the partition sum. To what extent they depend on the Hamiltonian is an interesting question.
VII Discussion
@@
We conclude our presentation by flagging several research directions motivated by our results. Our approach can be extended to models whose associated Liouvillian dynamics are expressed as sign problem-free sums. The quantum XY model is a notable example [39], which crucially admits a global $\text{U}(1)$ symmetry and hence can provide access to transport properties of conserved charges [48, 49]. The presence of disorder is also compatible with our method, it will be interesting to investigate signatures of Many-Body Localization in the evolution of operator strings. Another important direction is extracting real frequency dynamical spectral functions away from the high-frequency limit, as can, in principle, be via Lanczos coefficients expressed in terms of moment ratios [50, 51, 52]. We note that the latter might suffer from high sensitivity to statistical noise and may require an improvement on current algorithms. Nevertheless, seeing how our Monte Carlo method fares against the existing techniques [53, 54, 55] will be an interesting line of enquiry. Lastly, we expect the support size scaling functions to be universal and valid for a generic class of chaotic Hamiltonians, and predicting them theoretically is an intriguing open question. We leave the aforementioned possibilities to future efforts.
Acknowledgements.
@@
We thank Ehud Altman, Assa Auerbach, Thomas Scaffidi and Daniel Parker for fruitful discussions and comments. S.G. acknowledges support from the Israel Science Foundation (ISF) Grant no. 586/22 and the US–Israel Binational Science Foundation (BSF) Grant no. 2020264. U.B. acknowledges support from the Israel Academy of Sciences and Humanities through the Excellence Fellowship for International Postdoctoral Researchers. Computational resources were provided by the Intel Labs Academic Compute Environment and the Fritz Haber Center for Molecular Dynamics, Hebrew University. X.C. acknowledges support from a France-Berkeley Fund grant (Project #24_2023).
Appendix A The Monte Carlo Algorithm
@@
In this section, we outline our Markov chain Monte Carlo (MCMC) sampling algorithm. We first discuss the re-weighting procedure used in Eq. 8. The weights $\{w_{n}\}$ may be initialized as $w_{n}=1/(J^{2n}(2n)!)$. This choice is useful as it is an upper bound to the radius of converge of the $\mu_{2n}$ sequence [39]. In practice one only needs to keep track of weight ratios between consecutive $n$ sectors $w_{n}/w_{n+1}$. The reweighting procedure works iteratively. At each step, we take $N$ samples and determine the number of times $c_{n}$ that each sector is visited by the MCMC. We define $c_{n}$ as:
$c_{n}=\sum_{i=1}^{N}\delta_{n,n_{i}}\,.$ (15)
@@
A.1 Add-drop
Figure 6: Illustration of the add-drop moves with (a) $\mathcal{L}_{Z}$ Liouvillians (b) $\mathcal{L}_{X}$ Liouvillians (c) $\mathcal{L}_{ZZ}$ Liouvillians. The temporal direction is vertical.
The add-drop move is designed to allow transitioning between the different $n$ sectors by adding and removing a pair of identical local Liouvillians. As illustrated in Fig. 6, the move is defined as follows:
@@ A.1 Add-drop
We now determine the transition probabilities. Let $i$ be a configuration in the $n-1$ sector with the $(k-1)$-th and $k$-th Liouvillian slots unoccupied. Let $f$ be a configuration in the $n$ sector that is identical to $i$ except with these slots occupied at site $r$ by Liouvillians $\mathcal{L}_{r_{k-1},a_{k-1}}=\mathcal{L}_{r_{k},a_{k}}=:\mathcal{L}_{r,a}$. The two configurations are thus related by an add-drop move. Based on Eq. 8, detailed balance imposes the constraint:
$\frac{T_{i\rightarrow f}A_{i\rightarrow f}}{T_{f\rightarrow i}A_{f\rightarrow i}}=\frac{P_{f}}{P_{i}}$
$=\frac{w_{n}(O_{k-1}|\mathcal{L}_{r,a}|O_{k})(O_{k}|\mathcal{L}_{r,a}|O_{k+1})(2n)(2n-1)}{w_{n-1}(2n_{\max}-2n+2)(2n_{\max}-2n+1)}\,.$ (16)
Here, $P_{i},P_{f}$ are associated probabilities for states $i,f$; $T_{i\rightarrow f},T_{f\rightarrow i}$ are probabilities to propose the add/drop transitions and $A_{i\rightarrow f},A_{f\rightarrow i}$ are the acceptance probabilities for the moves. Now, proposing an add operation involves choosing a time slice, followed by choosing a site and then choosing a Liouvillian type; proposing a drop amounts to simply choosing a time slice. Therefore the proposition probabilities are
$T_{i\rightarrow f}=\frac{1}{2n_{\max}+1}\times\frac{1}{L^{d}}\times\frac{(O_{k-1}|\mathcal{L}_{r,a}|O_{k})(O_{k}|\mathcal{L}_{r,a}|O_{k+1})}{\sigma}\,,\,$
$T_{f\rightarrow i}=\frac{1}{2n_{\max}+1}\,.$ (17)
@@
$A_{i\rightarrow f}=\text{min}\left(1,\frac{w_{n}L^{d}\sigma(2n)(2n-1)}{w_{n-1}(2n_{\max}-2n+2)(2n_{\max}-2n+1)}\right)\,,\,$
$A_{f\rightarrow i}=\text{min}\left(1,\frac{w_{n-1}(2n_{\max}-2n+2)(2n_{\max}-2n+1)}{w_{n}L^{d}\sigma(2n)(2n-1)}\right)\,.$ (18)
A.2 Swap
Figure 7: Illustration of the swap move. We show two examples of Liouvillians that commute with each other: (a) $\mathcal{L}_{Z}$ and $\mathcal{L}_{ZZ}$ Liouvillians sharing a site. (b) Two $\mathcal{L}_{ZZ}$ Liouvillians sharing just one site. In addition to (a) and (b), identical Liouvillians (those of the same kind, on the same site; and along the same axis in case of $\mathcal{L}_{ZZ}$) and those on different sites with no overlap, trivially commute.
The swap move permutes through different configurations in the same $n$ sector using the commutation property of Liouvillian pairs, see Fig. 7 for an illustration. Before discussing the steps, we first define the notion of “ceiling” and “floor”. The ceiling (floor) of a Liouvillian is the temporally closest Liouvillian above (below, respectively) it which doesn’t commute with it. If a non-commuting Liouvillian is lacking, the ceiling/floor is then defined to be the initial/final temporal boundary. Now, the swap steps are as follows: